OpenStudy (becca9898):

What are the approximate solutions of 4x^2 + 3 = -12x to the nearest hundredth?

OpenStudy (campbell_st):

well rewriting the equation you have \[4x^2 + 12x + 3 = 0\] now use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] where a = 4, b = 12 and c = 3 and solve for x

OpenStudy (campbell_st):

then when you get your answers, round to 2 decimal places.